OpenStudy (anonymous):

HELP PLEASE!!? The length of an object's shadow varies directly with its height. A boy who is 4 feet 6 inches tall casts a shadow that is 6 feet 3 inches long. What is the height of a tree that casts a shadow that is 25 feet long? A. 18 B. 20 C. 16 D. 216

OpenStudy (anonymous):

draw a picture

OpenStudy (anonymous):

\[\frac{ 54}{ 75}= \frac{ x }{ 300} =\\\frac{ \cancel {(3) }}{ \cancel {(3) }} \frac{ 18 }{ 25}= \frac{ x }{ 300}\\\\\frac{ 18 }{ 25}= \frac{ x }{ 300}\\ 300* \frac{ 18 }{ 25}= x \\\\ 12 * 18= x \\\\ x=216\]